Researchers observed a sample of 152 meaningful interactions, of which 72 were led by children


A study examined the interactions in a children's museum (Interactions by visitors to the museum included show-and-tell, learning, teaching, refocusing participatory play, advocating, and disciplining interactions.) Researchers observed a sample of 152 meaningful interactions, of which 72 were led by children and BO were led by adult caregivers. Complete parts a through d below. 


a. Give a point estimate of the true proportion of all meaningful interactions in a children's museum that are led by children 

The point estimate is _______ (Type an integer or decimal rounded to three decimal places as needed.) 
Answer #1

n = 152

x =72

Point estimate = sample proportion = р = x / n = 72/152 = 0.4737

The Point estimate is 0.474